The rubber cord of a catapult has a cross-sectional area of 1.0 mm^2 and a total upstretched length 10.0 cm. It is stretched to a 120 cm and then released to project a missile of mass 50 g. From energy consideration or otherwise, calculate the velocity of projection, taking Young’s modulus for the rubber as 50 x 108 Nm^-2. State the assumptions made in your calculation.
1
Expert's answer
2020-10-29T07:03:06-0400

Since the potential tensile energy is "E_p=\frac {k\times (\Delta l) ^ 2} {2}" , where "k" is the coefficient of elasticity ,"\Delta l" is the change in body length. "k =\frac {E\times S} {l}" , where "l" length, "E" is Young's modulus, "S" is the cross-sectional area of the body.

According to the law of conservation of energy "E_p=E_c" , where "E_c" is the kinetic energy of the body, therefore

"\ frac {\frac {E\times S} {l }\times (\Delta l) ^ 2} {2} =\frac {m\times v ^ 2} {2}" , where "m" body weight, "v" body speed

"v= \sqrt{\frac{\frac{E\times S}{l}\times (\Delta l)^2}{m}} =\sqrt{\frac{\frac{50\times 10^8\times 10^{-6}}{0.1}\times 1.1^2}{0.05}}=1100" m/s.